# TabiBERT:A Large-Scale ModernBERT Foundation Model and Unified Benchmarking Framework for Turkish

### v1

@@ 4.4.1 Motivation and Design Principles
    Quality over quantity: All datasets undergo systematic review for annotation consistency, label reliability, and task validity
  2. 2.
@@
    Standardization: Unified train-validation-test splits with explicit splitting methodology for reproducibility
  3. 3.
@@
    Task diversity: Coverage of traditional NLU tasks alongside modern evaluation paradigms (retrieval, code understanding, domain specialization)
4.4.2 Dataset Selection and Quality Assessment
@@
We systematically review existing Turkish fine-tuning datasets and analyze both dataset quality and task relevance for assessing model capabilities. Datasets are selected based on three mandatory criteria:
  1. 1.
@@
This manual quality assessment ensures our evaluation benchmarks reflect realistic and meaningful downstream applications rather than artifacts of poor annotation or ambiguous task definitions.
4.4.3 Standardized Split Generation Methodology
A significant challenge in Turkish NLP evaluation is the absence of standardized dataset splits. Many existing datasets lack explicit train-validation-test divisions, leading researchers to create inconsistent splits that prevent fair cross-study comparisons. We address this through a systematic split generation methodology applied uniformly across all datasets:
@@ 4.4.3 Standardized Split Generation Methodology
  • For datasets with a train-test split only, we create a validation split from the training set, sized to match the test set, and apply this across all models.
@@
  • In cases with a validation-test split, we split validation into train and validation sets in 80% and 20% proportions, respectively.
  • When only a single combined split is present, we partition the data into train, validation, and test sets in 70%, 15%, and 15% proportions, respectively.
All splits use stratified sampling to preserve class distributions and ensure statistical validity. To support reproducibility and community adoption, we release these standardized evaluation suites publicly on HuggingFace (TabiBench).
TabiBench encompasses diverse task categories, combining curated existing Turkish datasets with newly created resources.
4.4.4 Task Categories and Dataset Coverage
Our fine-tuning and evaluation benchmark covers the following task categories and corresponding Turkish datasets:
@@ 4.4.4 Task Categories and Dataset Coverage
  • Natural language inference tasks include: SNLI-TR and MultiNLI-TR (BudurEtAl2020Data), which are derived from the Natural Language Inference in Turkish (NLI-TR) dataset.
  • Question answering tasks include: TQuAD (ErdometoTQuad2) and XQuAD (Artetxe2019CrossLingual), which provide Turkish datasets for reading comprehension and cross-lingual question answering.
  • Information retrieval tasks follow the Turkish translations provided by the TR-MTEB initiative (BaysanGungor2025TR-MTEB) and include: BiText, MsMarco-TR, Scifact-TR, Fiqa-TR, NFCorpus-TR, and Quora-TR. These datasets cover diverse domains such as general web search, scientific literature, financial documents, and question–answer retrieval.
@@
    • Apps-Retrieval (Hendrycks2021Measuring) consists of programming problem descriptions as queries and corresponding code solutions as documents. Its Turkish variant, Apps-TR, is created by translating only the problem descriptions.
@@
  • Academic domain tasks assess model performance on specialized scholarly text. Three datasets are translated from English using GPT-4.1, while one dataset is natively constructed in Turkish:
@@
    • Thesis-Abstract-Classification-11K is a natively constructed Turkish dataset derived from thesis abstracts collected from the National Thesis Center of Turkey (TurkishThesis2025Ertugrul). The task is to predict the field of study of a thesis abstract among 187 categories. A balanced subset with 60 samples per category is constructed, making this a challenging multi-class classification task due to the large label space and specialized academic language.
@@ 5 Evaluation
This section presents a comprehensive evaluation of TabiBERT across a diverse range of Turkish NLP tasks from TabiBench. We first describe the models compared in our evaluation, including state-of-the-art monolingual Turkish baselines and the multilingual mmBERT model. We then detail our evaluation methodology, including task and dataset selection, hyperparameter tuning procedures, fine-tuning configurations, context length selection, training objectives, and evaluation metrics. Finally, we present detailed results across eight task categories, demonstrating TabiBERT’s performance relative to baseline models and analyzing key strengths and limitations.
We begin by summarizing the key results:
  • TabiBERT achieves state-of-the-art performance among Turkish models, outperforming all monolingual baselines with a total average score of $77.58$, surpassing the previous best model by $1.62$ points.
  • The largest improvement is observed in question answering tasks, where TabiBERT achieves an F1 score of $69.71$, outperforming the next best Turkish model by $9.55$ points (1616% relative improvement).
  • TabiBERT leads the Turkish models in five out of eight task categories, including code retrieval where it outperforms the previous best by $2.41$ points, and information retrieval where it achieves an NDCG@10 score of $75.44$.
5.1 Models Compared
We compare TabiBERT with three Turkish BERT models, configurations of which are based on BERT-Base, representing the monolingual state-of-the-art approaches for Turkish. They are BERTurk, YTU-Cosmos-BERT, and TurkishBERTweet. YTU-Cosmos-BERT and TurkishBERTweet are selected on the fact that they have a clear recipe in the form of publications. Even though BERTurk is not the product of an academic work, it is a well-established and widely used model in Turkish NLP research as a strong baseline. Moreover, all of the mentioned models have a large, diverse pre-training corpus. In addition to Turkish BERT models, we also compare against the recently introduced ModernBERT-based multilingual encoder mmBERT.
BERTurk
(Schweter2020BERTurk) is a widely used Turkish monolingual encoder-only model, pre-trained on the Web and special corpora. It serves as a strong baseline for Turkish NLP tasks and has influenced much of the subsequent Turkish NLP research.
YTU-Cosmos-BERT
(Kesgin2023YTU) is an uncased Turkish BERT model, pre-trained on a large Turkish corpus including The Web, Wikipedia and books.
TurkishBERTweet
(Najafi2023TurkishBERTweet) TurkishBERTweet is another uncased Turkish BERT model, pre-trained in large scale, tailored for social media.
mmBERT
(Marone2025mmBERT) mmBERT is a multilingual encoder based on ModernBERT, with the aim of obtaining a multilingual ModernBERT model.
5.2 Methodology
We conduct a comprehensive evaluation of 5 models across various Turkish NLP tasks, combined under the TabiBench benchmark. To ensure fair and reproducible comparisons, we establish an end-to-end pipeline for task selection, dataset preparation, hyperparameter tuning, and fine-tuning.
Task and Dataset Selection
Our evaluation includes a diverse set of tasks that assess model capabilities across various domains and applications. These tasks include natural language understanding tasks such as text classification, natural language inference, semantic textual similarity, named entity recognition, and part-of-speech tagging, as well as more specialized tasks, including question answering, information retrieval, code retrieval, and academic domain tasks. We develop TabiBench, a comprehensive benchmark consisting of twenty eight datasets spanning these eight task types, as further elaborated in Section 4.4.
@@ 5.2 Methodology
We perform systematic hyperparameter tuning for all model-task pairs to ensure optimal performance. We determine our hyperparameter search space based on prior work (He2023DeBERTaV3; Warner2024ModernBERT; Breton2025NeoBERT) and our compute budget, ensuring effective fine-tuning across all tasks and models. Our hyperparameter search space is presented in Table 3.
@@
Early Stopping
All models are trained for up to 10 epochs with early stopping, with a patience of 2 epochs, based on validation split loss to prevent overfitting. The best hyperparameter configuration for each model-task pair is determined through validation set performance, and these configurations are then used for final fine-tuning and test set evaluation. For the uncased YTU-Cosmos-BERT and TurkishBERTweet models, lower-case conversion is applied to avoid problems with the Turkish character "I":
text = text.replace("I", "ı").lower()
@@
For each model-task pair, we fine-tune using the optimal hyperparameters identified during tuning. All fine-tuning experiments use a linear decay with warmup scheduler, where the learning rate warms up to the full value over 6% of the training duration and linearly decays to 0.02x the full learning rate by the end of training. We employ mixed precision training and set a random seed of 25 for reproducibility across all experiments. These configurations are based on the example finetuning configs released in the ModernBERT repository 333github.com/AnswerDotAI/ModernBERT/yamls/finetuning/hf-bert-base-uncased.yaml. For each fine-tuning task category, we report a single score by computing a weighted average across all datasets in that category, where each dataset’s weight is proportional to its test set size. This methodology ensures that larger, more representative datasets have a corresponding influence on the overall results, due to the variation in test set sizes ranging from 150 to $35,000$ examples.
@@
To determine appropriate context lengths for fine-tuning, we compute the 95th percentile of sequence lengths for each dataset. For each model-task pair, we set the maximum context length to the minimum of the model’s context length and the dataset’s 95th percentile length, making sure that no sequences are truncated unnecessarily while respecting each model’s architectural constraints. Since all other Turkish models have a context length of $512$ tokens, whereas TabiBERT and mmBERT support up to $8,192$ tokens, the latter two are able to utilize longer context windows for tasks where the 95th percentile exceeds $512$ tokens, providing them with an advantage on longer sequences.
@@
We apply task-specific training objectives for fine-tuning across different task types. For text classification and natural language inference tasks, we use cross-entropy loss, implemented through HuggingFace’s default loss computation for sequence classification models. Token classification tasks, including named entity recognition and part-of-speech tagging, are trained using cross-entropy loss applied at the token level. Semantic textual similarity tasks are formulated as regression problems, using a custom regression model wrapper that optimizes mean squared error (MSE) loss to predict continuous similarity scores. Question answering tasks employ span prediction, where the model predicts the start and end positions of answer spans within the context using cross-entropy loss for each position. For information retrieval and code retrieval tasks, we use contrastive learning objectives to learn representations that maximize the similarity between queries and relevant documents while minimizing similarity with irrelevant documents.
@@
We use task-appropriate evaluation metrics: macro-averaged F1 for classification and NLI (to account for class imbalance), micro-averaged F1 at the word level for token classification, Pearson correlation for semantic similarity, F1 for question answering, and NDCG@10 for information retrieval and code retrieval tasks.
@@ 5.3 Results
The performance of TabiBERT compared to monolingual Turkish models and mmBERT across all tasks is presented in Table 4. TabiBERT outperforms all other Turkish models in total average score across all tasks, achieving a score of 77.58, surpassing the previous best Turkish model, BERTurk, by 1.62 points. Moreover, the improvement is consistent across tasks with TabiBERT achieving the highest score in 5 out of 8 task categories, and being competitive in the remaining tasks. Despite being optimized for long context, TabiBERT maintains strong performance across tasks, demonstrating its versatility and effectiveness across a wide range of applications.
@@
Model # of params Text Clf Token Clf STS NLI QA Academic Retrieval Code Retrieval Total Avg
@@
Core Natural Language Understanding Tasks
Being always either on par with or slightly above the strongest Turkish baselines, TabiBERT demonstrates strong performance across core natural language understanding tasks: text classification, token classification, semantic textual similarity, and natural language inference. Text classification results of all models except for TurkishBERTweet are relatively close to each other, with YTU-Cosmos-BERT achieving the highest score of 84.25, followed closely by TabiBERT with 83.44. All Turkish models achieve high scores above 92 in token classification tasks with BERTurk leading with 93.67, closely followed by YTU-Cosmos-BERT and TabiBERT with 93.60 and 93.42, respectively. As it is in the case of text classification, all models except for TurkishBERTweet achieve similar scores in semantic textual similarity tasks with BERTurk leading with 85.33, followed by TabiBERT with 84.74. TabiBERT takes the lead in natural language inference tasks with a score of 84.51, slightly surpassing BERTurk’s score of 84.33 and YTU-Cosmos-BERT’s score of 84.16. In STS and NLI tasks, TabiBERT consistently matches or marginally exceeds BERTurk and YTU-Cosmos-BERT, demonstrating that the native long context of TabiBERT does not come at the expense of sentence-level semantic precision. The close scores across the top three models indicate that these tasks are close to saturation.
Question Answering
The largest margin of improvement for TabiBERT over other Turkish models is observed in question answering tasks. TabiBERT outperforms all monolingual baselines by a considerable margin, achieving an F1 score of 69.71 compared to the next best score of 60.16 by BERTurk. There are two factors in explaining this large margin: 1) TabiBERT excels at more difficult tasks. 2) TabiBERT’s long context length allows it to process long contexts in the TQuAD dataset without information loss to truncation.
Academic Domain Tasks
Performance gains are also evident in specialized academic tasks, including NLI and multi-class text classification. Trained on a corpus enriched with Turkish scientific articles and graduate theses, TabiBERT achieves the highest score among Turkish models, with particularly strong results on Thesis-Abstract-Classification-11K, a challenging long context multi-class task with limited per-class data. Thus, TabiBERT’s pre-training on academic texts and its long context capabilities contribute to its superior performance in this domain.
Retrieval and Code Retrieval
TabiBERT outperforms all other Turkish models in information retrieval and code retrieval tasks. TabiBERT consistently surpasses prior monolingual models on TRMTEB retrieval benchmarks, achieving an NDCG@10 score of 75.44 and surpassing the 74.84 score of the previous best Turkish model, BERTurk. The improvement is even more evident in code retrieval tasks, where TabiBERT achieves an NDCG@10 score of 56.95, considerably outperforming BERTurk by 2.41 points and YTU-Cosmos-BERT by 3.15 points. The benefits of extended context length and exposure to heterogeneous pre-training data, including bilingual text and code snippets, are particularly pronounced in retrieval and code retrieval tasks.
Comparison with Multilingual mmBERT
While TabiBERT outperforms all monolingual Turkish models, it still lags behind the multilingual mmBERT model, which achieves a total average score of 79.26. The explanation is threefold:
  • mmBERT is a larger model with 307307M parameters, compared to TabiBERT with 149149M parameters.
  • mmBERT’s tokenizer spends 41% more tokens than TabiBERT’s tokenizer to represent Turkish texts.
  • mmBERT is pre-trained on a massive multilingual corpus for a longer duration.
Although mmBERT still follows the BERT-base configuration with the same number of non-embedding parameters, despite its higher memory requirements, its large embedding matrix provides it with greater capacity to learn complex patterns. Moreover, mmBERT’s 41% higher fertility 4.1 results in greater fragmentation of Turkish texts, allowing it to better capture the nuances of Turkish morphology, as the tokenization process breaks words into smaller character ngrams. Although this comes at a cost, it allows mmBERT to perform better by allocating more compute on a given text. mmBERT’s pre-training on a massive multilingual corpus of $1,800$ languages, including high-resource languages such as English, enables it to learn richer representations by leveraging knowledge across languages. Not only that, but it is also pre-trained for a total of 33T tokens, significantly more than TabiBERT’s 11T tokens, which has been only possible by conducting more than $11.83$ epochs over the relatively small Turkish corpus.
Despite its strong performance across Turkish NLP tasks, mmBERT’s large size and multilingual nature may introduce inefficiencies for applications focused solely on Turkish. As a result of its large embedding matrix and the total number of parameters, mmBERT has significantly higher memory requirements than TabiBERT. Moreover, although both models have a native context length of $8,192$ tokens, mmBERT tokenizer’s high fertility results in increased token counts for Turkish texts and a 29% reduction in effective context length compared to TabiBERT.
Comparison Against Task-Specific Best Models
TabiBERT remains competitive when compared against the best-performing model for each individual task, as shown in Table 5, It achieves an average score of $77.58$, surpassing the task-specific best models’ average of $76.11$ by $1.47$ points. Although this improvement is slightlysmaller than in the comparison against all models, TabiBERT still outperforms the task-specific best in 5 of the 8 task categories, with particularly strong improvements in question answering ($9.55$ points), code retrieval ($2.41$ points), and academic domain tasks ($0.66$ points). For the remaining tasks where TabiBERT does not exceed the task-specific best, the performance differences are minimal, ranging from 0.81-0.81 to 0.25-0.25 points, indicating that TabiBERT remains highly competitive even against the strongest models for individual tasks.
Table 5: TabiBERT performance compared to the previous best-performing Turkish model on each downstream task. For each task, we report TabiBERT’s score, the previous best model’s score and name, and the absolute difference ($\Delta$). Boldfaced values indicate the best score per task.

### v3

@@ 4 TabiBench: A Standardized Turkish NLP Benchmark
  • Quality over quantity: Systematic review of datasets for annotation consistency, label reliability, and task validity.
@@
  • Standardization: Fixed train-validation-test splits with explicit methodology for reproducibility and fair comparison.
@@
  • Task diversity: Inclusion of modern NLP tasks such as retrieval, code understanding, and domain specialization, as well as traditional NLU tasks.
By addressing the fragmentation and quality concerns evident in prior Turkish NLP benchmarks, TabiBench aims to enable more reliable, comprehensive, and future-facing evaluation of Turkish encoder models. The remainder of this section describes the task categories and dataset coverage in detail.
4.1 Dataset Selection and Quality Assessment
@@
We systematically review existing Turkish fine-tuning datasets, evaluating both dataset quality and task relevance. To ensure our evaluation benchmark reflects realistic and meaningful downstream applications, datasets are selected based on three criteria:
@@
4.2 Standardized Split Generation
Another issue encountered in Turkish NLP evaluation is the absence of standardized dataset splits. Many existing datasets lack explicit train-validation-test divisions, leading researchers to create inconsistent splits that prevent fair cross-study comparisons. We address this through a systematic split generation methodology applied uniformly across all datasets:
@@ 4.2 Standardized Split Generation
  • For datasets with a train-test split only, we create a validation split from the training set, sized to match the test set.
@@
  • In cases with a validation-test split, we split validation into train and validation sets in 80 and 20 per cent proportions, respectively.
  • When only a single combined split is present, we partition the data into train, validation, and test sets in 70, 15, and 15 per cent proportions, respectively.
All splits use stratified sampling to preserve class distributions and ensure statistical validity.
4.3 Task Categories and Dataset Coverage
TabiBench encompasses 28 datasets spanning 8 major task categories: text classification (4 datasets), token classification (4 datasets), semantic textual similarity (2 datasets), natural language inference (2 datasets), question answering (2 datasets), information retrieval (6 datasets), code retrieval (4 datasets), and academic understanding tasks (4 datasets). Our benchmark combines 21 carefully curated Turkish datasets with 7 newly created machine-translated datasets for code retrieval and academic understanding evaluation, ensuring both depth and breadth in evaluation. The following sections detail each task category and its corresponding datasets:
@@ 4.3 Task Categories and Dataset Coverage
  • Natural language inference tasks include: SNLI-TR and MultiNLI-TR, which are derived from the Natural Language Inference in Turkish (NLI-TR) (BudurEtAl2020Data) dataset.
  • Question answering tasks include: TQuAD (ErdometoTQuad2) and XQuAD (Artetxe2019CrossLingual), for reading comprehension and cross-lingual question answering.
  • Information retrieval tasks include TR-MTEB (BaysanGungor2025TR-MTEB), obtained by translating the original MTEB (Muennighoff2023MTEB): BiText, MsMarco-TR, Scifact-TR, Fiqa-TR, NFCorpus-TR, and Quora-TR, covering diverse domains from general web search to scientific and financial documents.
@@
    • AppsRetrieval (Hendrycks2021Measuring) consists of programming problem descriptions as queries and corresponding code solutions as documents. Its Turkish variant, Apps-TR, is created by translating only the problem descriptions.
@@
  • Academic understanding tasks assess model performance on specialized scholarly text. Three datasets are translated from English using GPT-4.1, while one dataset is natively constructed in Turkish:
@@
    • Thesis-Abstract-Classification-11K is a natively constructed Turkish dataset derived from thesis abstracts collected from the National Thesis Center of Turkey (TurkishThesis2025Ertugrul). The task is to predict the field of study of a thesis abstract among 187 categories. A balanced subset with 60 samples per category is constructed, making this a challenging multi-class classification task due to the large label space, small number of samples per category, and specialized academic language.
4.4 Impact and Availability
TabiBench addresses evaluation infrastructure gaps common to many lower-resource languages lacking standardized benchmarks comparable to GLUE (Wang2018GLUE) and SuperGLUE (Wang2019SuperGLUE) for English. We believe that this approach may serve as a valuable template to be applied to other low/medium resource languages facing similar challenges.
To support reproducibility and community adoption, we release TabiBench on HuggingFace (TabiBench) to: (1) establish evaluation best practices for Turkish NLP, (2) enable direct cross-model comparisons for future research, and (3) provide reproducible infrastructure for measuring progress. Beyond supporting this work, TabiBench represents an independent contribution that can facilitate more rigorous Turkish NLP research.
@@ 5 Evaluation
This section presents a comprehensive evaluation of TabiBERT across diverse Turkish NLP tasks. We first introduce the baseline models used for comparison, then describe our evaluation methodology, and present results demonstrating TabiBERT’s strengths and limitations. Prior to getting into the details, we can note the key findings:
  • TabiBERT achieves state-of-the-art performance among monolingual Turkish models with an average score of 77.58, surpassing the previous best by 1.62 points.
  • It leads in 5 of 8 task categories, with the largest gains in question answering (+9.55 points, 16 per cent relative improvement), code retrieval (+2.41 points, 4.4 per cent relative improvement).
  • Performance on text understanding tasks (classification, NER, STS, NLI) is competitive with established models, demonstrating that long-context optimization does not compromise short-text understanding.
5.1 Models Used for Comparison
We compare TabiBERT against three monolingual Turkish encoder-only models based on BERT-Base configuration: BERTurk (Schweter2020BERTurk), ytu-cosmos-BERT (Kesgin2023YTU), and TurkishBERTweet (Najafi2023TurkishBERTweet). BERTurk, though not peer-reviewed, is long-standing and widely adopted in Turkish NLP; ytu-cosmos-BERT and TurkishBERTweet are peer-reviewed and well-documented. All three are pretrained on large, diverse Turkish corpora, constituting strong representative baselines.
[…]
5.2 Evaluation Methodology
@@
We perform a systematic hyperparameter search for all model-task pairs to ensure fair comparisons. The search space, informed by prior work (He2023DeBERTaV3; Warner2024ModernBERT; Breton2025NeoBERT) is presented in Table 3. All models train for up to 10 epochs with early stopping (patience=2) based on validation loss. Best configurations are selected via validation performance for the final test evaluation.
@@
[…]
Fair comparison protocol
Note that we perform hyperparameter tuning followed by fine-tuning for all models in our evaluation using TabiBench, not just TabiBERT. This ensures all models are evaluated at their best performance on these tasks, making comparisons attributable to architectural differences and explicit design decisions rather than hyperparameter selection discrepancies.
@@
We compute the 95th percentile sequence length for each dataset and set the maximum context length to $\min(\text{model context limit},\text{dataset 95th percentile})$. If the 95th percentile exceeds the model’s maximum context length, we use the model limit. This design choice enables efficient use of available computer resources while maximizing context utilization. Previous Turkish BERT models are limited to 512 tokens, whereas TabiBERT and mmBERT support input lengths of up to $8,192$ tokens, enabling the use of longer contexts when available.
@@
Task-specific objectives are used across the benchmark: cross-entropy is utilized for text classification, NLI, and token classification; Mean Squared Error (MSE) loss is applied for semantic similarity regression; span prediction via cross-entropy is implemented for question answering; and contrastive learning is used for retrieval tasks.
Uncased models
For uncased models (ytu-cosmos-BERT, TurkishBERTweet), we apply Turkish-aware lowercasing (text.replace("I", "ı").lower()).
@@
Task-appropriate metrics are adopted for performance assessment: macro-averaged F1 is used for text classification and NLI; micro-averaged F1 at the word level is used for token classification; the Pearson correlation coefficient is used for semantic textual similarity (STS); the F1 score is used for question answering; and Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) is implemented for retrieval tasks.
Aggregation of results
For each task category, we report a single score via weighted average across datasets, with weights proportional to test set size (ranging 150 to 35K examples). This ensures that the aggregate score reflects performance across all test instances rather than giving equal weight to small and large datasets, which would overrepresent performance on smaller datasets. The detailed results per dataset for each task category are presented in Appendix A.
@@ 5.3 Results
TurkishBERTweet(163M)YTU-CosmosBERT( [^1] M)BERTurk(110M)TabiBERT(149M)mmBERT(307M)0202040406060808010010067.4867.4872.2672.2675.9675.9677.5877.5879.2679.26Total Avg. Score
(a) Overall average performances across all tasks on models. Values in parentheses denote the number of parameters if model in millions.
Text ClfTok. ClfSTSNLIQAAcad. Und.Info Retr.Code Retr.02020404060608080100100Score (0–100)TurkishBERTweetYTU-Cosmos-BERTBERTurkTabiBERTmmBERT
(b) Average performances by tasks: Text&NLI&Academic (macro-F1), Token Clf (micro-F1), STS (Pearson), QA (F1), and Information Retrieval&Code Retrieval (NDCG@10). TabiBERT consistently performs strongly across all tasks.
Figure 2: Overall and task-specific performance of models on TabiBENCH.
Table 4 and Figure 2 present comprehensive results across all models and tasks. TabiBERT achieves an overall average of 77.58, surpassing the previous best Turkish model (BERTurk, 75.96) by 1.62 points. Performance is consistent across categories, with TabiBERT leading in 5 of 8 tasks and remaining competitive in others.
@@
Model # of params Text Clf Token Clf STS NLI QA Academic Understanding Information Retrieval Code Retrieval Total Avg
@@
5.3.1 Overall Performance
[…]
Table 5 compares TabiBERT against the top-performing baseline for each task. TabiBERT achieves a mean score of 77.58, surpassing the aggregate best-performer average of 76.11. The model establishes new state-of-the-art results on five of the eight benchmark tasks. In these instances, TabiBERT outperforms the closest competitor by an average of 2.63 points, with significant gains in QA (+9.55 points) and Code Retrieval (+2.41 points) over BERTurk. Conversely, in cases where TabiBERT trails, the performance gap remains marginal, with an average deficit of 0.55 points and a maximum of 0.81 points.
These results suggest that TabiBERT’s architectural advantages are task-specific rather than uniform. Specifically, the model’s dominance in QA and retrieval tasks likely reflects the benefits of the ModernBERT extended context window, which facilitates reasoning over longer text spans. While further evaluation across broader long-context benchmarks is required to confirm this mechanism, the current findings indicate that the architectural advancements of ModernBERT translate effectively to the Turkish language domain.
Table 5: Performance difference ($\Delta$) between TabiBERT and previous best Turkish model across benchmark tasks. Tasks are ordered by relative model performance differences, ranging from the largest negative (trailing) to the largest positive (leading) gap. Positive values indicate tasks where TabiBERT achieves state-of-the-art results; negative values indicate it trails. TabiBERT leads on 5/8 tasks with an average improvement of +2.63 points and trails on 3/8 tasks by an average of 0.55 points.